# Deep Bayes Factors

### v2

@@ Abstract
The is no other model or hypothesis verification tool in Bayesian statistics that is as widely used as the Bayes factor. We focus on generative models that are likelihood-free and, therefore, render the computation of Bayes factors (marginal likelihood ratios) far from obvious. We propose a deep learning estimator of the Bayes factor based on simulated data from two competing models using the likelihood ratio trick. This estimator is devoid of summary statistics and obviates some of the difficulties with ABC model choice. We establish sufficient conditions for consistency of our Deep Bayes Factor estimator as well as its consistency as a model selection tool. We investigate the performance of our estimator on various examples using a wide range of quality metrics related to estimation and model decision accuracy. After training, our deep learning approach enables rapid evaluations of the Bayes factor estimator at any fictional data arriving from either hypothesized model, not just the observed data $Y_{0}$. This allows us to inspect entire Bayes factor distributions under the two models and to quantify the relative location of the Bayes factor evaluated at $Y_{0}$ in light of these distributions. Such tail area evaluations are not possible for Bayes factor estimators tailored to $Y_{0}$. We find the performance of our Deep Bayes Factors competitive with existing MCMC techniques that require the knowledge of the likelihood function. We also consider variants for posterior or intrinsic Bayes factors estimation. We accompany our model selection framework with a model criticism framework based on predictive generative Bayes. We demonstrate the usefulness of our approach on a relatively high-dimensional real data example about determining cognitive biases.
Bayes Factors,
Likelihood-free Bayesian Inference,
Classification,
Implicit Models,
keywords:
volume: TBAissue: TBA
[…]
t1The author gratefully acknowledges support from the James S. Kemper Foundation Faculty Research Fund at the Booth School of Business as well as the National Science Foundation (DMS:1944740).
@@ 1 Introduction
Bayes factors (BFs) are the main staple of Bayesian hypothesis testing and model comparisons (Jeffreys, 1935; Berger, 1985; Kass and Raftery, 1995). Defined as ratios of two marginal likelihoods, Bayes factors augment the classical likelihood ratio test by reflecting relative predictive performance. In addition, they quantify the extent to which data warrant prior-to-posterior change in model odds.
@@
While Bayes factors have a wide advocacy within the Bayesian community Berger and Sellke (1987); Savage (1972), there are operational and conceptual connundra. Besides the difficulty in choosing plausible default proper priors for model comparisons (O’Hagan, 1995; Berger and Pericchi, 1996), Bayes factor computation itself can be daunting. Practitioners may resort to convenient approximations such as the BIC (Schwarz, 1978) which, unfortunately, may not easily lend itself to an analysis that seeks to take advantage of background knowledge through informative priors. Indeed, the BIC approximates the “default” Bayes factor that is based on a unit-information prior (Kass and Wasserman, 1995). When likelihoods are, in fact, available, various analytical approximations (Jeffreys, 1998; Kass and Raftery, 1995; Raftery, 1996) or simulation approaches (including Bridge sampling and variants thereof Meng and Schilling (2002); Gelman and Meng (1998)) are available. Our methodology targets situations when these approaches cannot be readily applied. We focus on the feasibility of Bayes factor computation in the context of simulation-based (generative) models. Bayes factors are used broadly across various disciplines (Geweke, 1999; Bridges et al., 2009; Minin et al., 2003; Lee and Wagenmakers, 2014) including machine learning where the use of marginal likelihoods for model comparisons is increasingly more appreciated (Qi and Zhang, 2001; Immer et al., 2021; Dutordoir et al., 2020). Our purpose is to broaden the reach of Bayes factors to contexts involving complex models where marginal likelihood evaluations would seem nonviable.
@@
A wide range of sampling-based Bayes factor estimation techniques exists, largely when the likelihood is accessible. Rubin (Rubin, 1987; DB, 1988) emphasized importance sampling for marginal likelihoods, noting however its potential instabilities when the posterior distribution is sharply peaked relative to the prior. Savage-Dickey representations, as laid out by Verdinelli and Wasserman (1995), cater specifically to nested models. Additional methods worth noting include harmonic means, Laplace approximation, stochastic substitution, nested sampling, and reverse logistic regression (Raftery et al., 2006; Tierney and Kadane, 1986; Gelfand and Smith, 1990; Geyer, 1994; Chopin and Robert, 2010). Comparative insights into these methodologies can be sourced from Marin et al. (2007). Our contribution centers primarily on situations when likelihoods are implicit (analytically intractable), not only when the computation of marginal likelihoods presents a challenge. A crucial requirement for our methodology is that the likelihood can be simulated from. Bayesian inference in such likelihood-free scenarios has been accessed through the utilization of Approximate Bayesian Computation (ABC) (Tavaré et al., 1997; Pritchard et al., 1999; Beaumont et al., 2002). A default ABC approach to estimating Bayes factors involves calculating the ratio of acceptance rates under considered models using ABC reference tables (Grelaud et al., 2009; Toni and Stumpf, 2010; Didelot et al., 2011). However, Robert et al. (2011) show that ABC for Bayes factor calculations can be biased when summary statistics are insufficient for model selection. We investigate the usefulness of deep learning for Bayes factor estimation. The core idea is fairly simple. We deploy the likelihood-ratio trick from machine learning (Hastie et al., 2009; Sugiyama et al., 2012; Mohamed and Lakshminarayanan, 2016) to estimate marginal likelihood ratios via binary classification. The effectiveness of classification-based density ratio estimates has been appreciated in the Bayesian context for Metropolis-Hastings simulation (Kaji and Ročková, 2022), posterior estimation (Thomas et al., 2022), and ABC (Wang et al., 2022). We were intrigued by the possibility of leveraging contrastive learning for Bayes factor computation in models where prior and likelihood simulations are possible.
@@
In this work, we introduce Deep Bayes Factor (DeepBF), a neural classifier trained on simulated datasets to learn a mapping whose functional constitutes a Bayes factor estimator. The trained mapping enables Bayes factor evaluations over the entire data domain (not just observed data) with negligible additional computational cost. This feature allows us to compute distributions of Bayes factors under each hypothesis and perform model selection inference using tail area evidence measures. Many Bayes factor computation methods are tailored to the observed data (Meng and Wong, 1996; Meng and Schilling, 2002) which allow only point-evaluations at observed data as opposed to inference using plausible data under each hypothesis. We support our development theoretically by providing sufficient conditions under which our Deep Bayes Factors are consistent estimators of Bayes factors as well as sufficient consistency conditions for model selection. Viewing the Bayes factor through the lens of binary classification aligns with Pudlo et al. (2016), who recast ABC model selection as a classification problem. They employ random forests to select a model by a majority vote. Instead, we focus on binary classification where the purpose is to learn marginal likelihood ratios. Contrary to the method in Pudlo et al. (2016), our strategy circumvents a secondary learning phase for gauging model posterior estimates, delivering results in only one stage. There have several attempts to apply deep learning to the problem of Bayes factor estimation from other perspectives. For example, Jia and Seljak (2020) used normalizing flow approach to estimate the marginal likelihood, and Xing (2022) suggested to use f-GAN to improve bridge estimators (Ali and Silvey, 1966). While their approaches rely on MCMC sampling from the posteriors to obtain training dataset, our DeepBF is trained solely on simulated data from two competing models, not the posteriors.
@@
Our second contribution is furnishing our methodology with a toolkit for evaluating goodness of model fit as well as goodness of Bayes factors for model selection inference. Our exhaustive evaluation toolkit provides a more comprehensive framework for informed model selection and enables effective objective comparisons of various Bayes factor estimators. Firstly, since model selection cannot be entirely separated from model criticism, we discuss several goodness-of-fit strategies using contrastive learning, which can be used as a prerequisite to Bayes factor evaluations. In particular, we develop a contrastive learning approach to model adequacy assessment based on comparing observed data with data generated from a posterior predictive distribution. We develop a score based on an estimated classifier and gauge model adequacy based its the departure away from $1/2$. While Bayes factor comparisons indicate which of the two considered models is more plausible, our model adequacy framework helps determine whether considered models themselves are compatible with the observed data. Secondly, we evaluate goodness of Bayes factors with visualization tools and validation measures including ROC curves and estimated model prior quality measures. Additionally, while the magnitude of Bayes factors evaluated at observed data traditionally indicates the strength of model evidence (Jeffreys, 1998; Kass and Raftery, 1995; Lee and Wagenmakers, 2014), recent works complement this by quantifying the amount of evidence using a tail probability (Baskurt and Evans, 2013) derived from simulated data (García-Donato and Chen, 2005; Gu et al., 2016; Schönbrodt and Wagenmakers, 2018; Schad et al., 2022). This improved understanding can lead to calibrations of Bayes factor decision thresholds. Our validation measures, therefore, include the distributional quality of Bayes factor estimates that does not penalize systematic bias or imprecise scale such as area under ROC curve (AUC) and MSE of surprise measures. These evaluations are further enhanced by our rapid and simultaneous computation of our DeepBF estimator. We demonstrate performance of Deep Bayes Factors on several well-studied examples where ABC model choice faces challenges. Lastly, we apply our approach to a real data example on testing for the presence of cognitive biases.
@@ 2 Approximating Bayes Factors via Classification
Our setup consists of an observed data vector $Y^{(n)}_{0}=(Y_{1}^{0},\dots,Y_{n}^{0})\in\mathcal{Y}^{n}$ which could have plausibly arrived from a model $M_{1}$ or a model $M_{2}$. These two models are specified through conditional densities $\pi_{1}(\cdot\,|\>\theta_{1})$ and $\pi_{2}(\cdot\,|\>\theta_{2})$ with unknown model parameters $\theta_{1}\in\mathbb{R}^{d_{1}}$ and $\theta_{2}\in\mathbb{R}^{d_{2}}$ that are assigned prior distributions $\pi_{1}(\theta_{1})$ and $\pi_{2}(\theta_{2})$, respectively. The Bayes Factor is then defined as a ratio of the marginal likelihoods evaluated at $Y^{(n)}_{0}$
@@
When likelihoods are analytically unavailable, but can be simulated from, a popular approach to estimating Bayes factors is ABC (Grelaud et al., 2009; Toni and Stumpf, 2010; Didelot et al., 2011; Robert et al., 2011). We review this approach to build a foundation for our own development.
@@ 2.1 ABC Model Choice
ABC was originally introduced in population genetics (Tavaré et al., 1997; Pritchard et al., 1999; Beaumont et al., 2002) for inference about evolution of species. Instead of sampling directly from the posterior, ABC draws candidate parameters from the prior and accepts them if simulated data from the likelihood (given candidate parameters) closely aligns with the observed data. For a comprehensive introduction to ABC we refer to Sunnåker et al. (2013). With some caution, ABC can be also used for model choice.
A common ABC approach to Bayes factor estimation (Robert et al., 2011) is calculating the ratio of acceptance rates for the two models using ABC reference tables involving simulated data from the marginal likelihood. During the $i^{th}$ of $N$ total ABC iterations, one samples the model index $k^{(i)}\in\{1,2\}$ based on prior model probabilities $\pi(M_{1})$ and $\pi(M_{2})$. Second, parameters $\theta^{(i)}$ are drawn from the prior $\pi_{k^{(i)}}(\cdot)$ and, finally, an $(n\times 1)$ fake data vector $Y^{(i)}$ is simulated from $\pi_{k^{(i)}}(\cdot\,|\>\theta^{(i)})$. The estimate of the Bayes factor based on the ABC lookup table $(k^{(i)},\theta^{(i)},Y^{(i)})_{i=1}^{N}$ is then obtained as
@@
The implementation of $\widehat{BF}^{\rm ABC}_{1,2}(Y_{0}^{(n)})$ in (2) is straightforward and bypasses direct calculations of the likelihood which makes it applicable to a wide range of complex and non-standard models. For Gibbs random fields, Grelaud et al. (2009) demonstrated convergence of (2) to the true Bayes factor. Robert et al. (2011) cautioned against inherent bias in ABC-based Bayes factor estimates due to the information loss induced by summary statistics. Their core intuition is that, even if the summary statistic $\eta(\cdot)$ is sufficient for both $\pi_{1}(\cdot\,|\>\theta_{1})$ and $\pi_{2}(\cdot\,|\>\theta_{2})$, it may not be sufficient for $\{(k,\pi_{k}(\cdot\,|\>\theta_{k}))\}_{k=1,2}$, rendering it unsuitable for model selection.
Previous studies in ABC literature have also highlighted the relationship between Bayes factor estimation and the classification problem. Given a set of $K$ competing models $M_{1},\cdots,M_{K}$, Fagundes et al. (2007) estimated the posterior probability $\pi(M_{k}\,|\>Y_{0}^{(n)})$ of each model using multinomial logistic regression trained on an ABC filtered reference table using the model category $k^{(i)}$ as a label and summarized data $\eta(Y^{(i)})$ as predictors. A similar concept is presented in Pudlo et al. (2016) who employed two distinct random forests in two stages. First, a random forest classifier is trained. In the second stage, a separate random forest regression is deployed, targeting the error rate of the initial classifier as an estimator of $\pi(M_{k}\,|\>Y^{(n)}_{0})$. Our method operates in a single stage, where a more flexible classifier is deployed and regarded as a function of the Bayes factor. While we also frame Bayes factor estimation as classification, we do not rely on summary statistics and acceptance thresholds.

### v3

@@ Abstract
There is no other model or hypothesis verification tool in Bayesian statistics that is as widely used as the Bayes factor. We focus on generative models that are likelihood-free and, therefore, render the computation of Bayes factors (marginal likelihood ratios) far from obvious. We propose a deep learning estimator of the Bayes factor based on simulated data from two competing models using the likelihood ratio trick. This estimator is devoid of summary statistics and obviates some of the difficulties with ABC model choice. We establish sufficient conditions for consistency of our Deep Bayes Factor estimator as well as its consistency as a model selection tool. We investigate the performance of our estimator on various examples using a wide range of quality metrics related to estimation and model decision accuracy. After training, our deep learning approach enables rapid evaluations of the Bayes factor estimator at any fictional data arriving from either hypothesized model, not just the observed data $Y_{0}$. This allows us to inspect entire Bayes factor distributions under the two models and to quantify the relative location of the Bayes factor evaluated at $Y_{0}$ in light of these distributions. Such tail area evaluations are not possible for Bayes factor estimators tailored to $Y_{0}$. We find the performance of our Deep Bayes Factors competitive with existing MCMC techniques that require the knowledge of the likelihood function. We also consider variants for posterior or intrinsic Bayes factors estimation. We accompany our model selection framework with a model criticism framework based on predictive generative Bayes. We demonstrate the usefulness of our approach on a relatively high-dimensional real data example about determining cognitive biases.
Abstract
This supplementary material contains proofs of all the theorems in the main text as well as the description of the rank-based ABC algorithm (in Section S2). In particular, the proof of the consistency of our Deep Bayes Factor estimator in Theorem 1 is presented in Section S1.2. The proofs for its consistency as a model selection tool in Theorem 2 is presented in Section S1.3. Practical considerations of choosing the deep learning architecture are elaborated in Section S3. Section S4 demonstrates promising results achieved by using DeepSets for Deep Bayes Factors. Section S5 provides experimental details, and Section S6 provides some additional empirical investigations, and Section S7 contains some additional visualizations and tables.
Keywords: Bayes Factors, Likelihood-free Bayesian Inference, Classification, Implicit Models
@@ 1 Introduction
Bayes factors (BFs) are the main staple of Bayesian hypothesis testing and model comparisons [37, 8, 43]. Defined as ratios of two marginal likelihoods, Bayes factors augment the classical likelihood ratio test by reflecting relative predictive performance. In addition, they quantify the extent to which data warrant prior-to-posterior change in model odds.
@@
While Bayes factors have a wide advocacy within the Bayesian community [10, 71], there are operational and conceptual connundra. Besides the difficulty in choosing plausible default proper priors for model comparisons [57, 9], Bayes factor computation itself can be daunting. Practitioners may resort to convenient approximations such as the BIC [75] which, unfortunately, may not easily lend itself to an analysis that seeks to take advantage of background knowledge through informative priors. Indeed, the BIC approximates the “default” Bayes factor that is based on a unit-information prior [44]. When likelihoods are, in fact, available, various analytical approximations [38, 43, 63] or simulation approaches (including Bridge sampling and variants thereof [51, 22]) are available. Our methodology targets situations when these approaches cannot be readily applied. We focus on the feasibility of Bayes factor computation in the context of simulation-based (generative) models. Bayes factors are used broadly across various disciplines [23, 14, 53, 47] including machine learning where the use of marginal likelihoods for model comparisons is increasingly more appreciated [62, 35, 17]. Our purpose is to broaden the reach of Bayes factors to contexts involving complex models where marginal likelihood evaluations would seem nonviable.
@@
A wide range of sampling-based Bayes factor estimation techniques exists, largely when the likelihood is accessible. Rubin [69, 70] emphasized importance sampling for marginal likelihoods, noting however its potential instabilities when the posterior distribution is sharply peaked relative to the prior. Savage-Dickey representations, as laid out by [85], cater specifically to nested models. Additional methods worth noting include harmonic means, Laplace approximation, stochastic substitution, nested sampling, and reverse logistic regression [64, 82, 20, 24, 15]. Comparative insights into these methodologies can be sourced from [49]. Our contribution centers primarily on situations when likelihoods are implicit (analytically intractable), not only when the computation of marginal likelihoods presents a challenge. A crucial requirement for our methodology is that the likelihood can be simulated from. Bayesian inference in such likelihood-free scenarios has been accessed through the utilization of Approximate Bayesian Computation (ABC) [80, 60, 5]. A default ABC approach to estimating Bayes factors involves calculating the ratio of acceptance rates under considered models using ABC reference tables [27, 84, 16]. However, [66] show that ABC for Bayes factor calculations can be biased when summary statistics are insufficient for model selection. We investigate the usefulness of deep learning for Bayes factor estimation. The core idea is fairly simple. We deploy the likelihood-ratio trick from machine learning [31, 78, 54] to estimate marginal likelihood ratios via binary classification. The effectiveness of classification-based density ratio estimates has been appreciated in the Bayesian context for Metropolis-Hastings simulation [41], posterior estimation [81], and ABC [90]. We were intrigued by the possibility of leveraging contrastive learning for Bayes factor computation in models where prior and likelihood simulations are possible.
@@
In this work, we introduce Deep Bayes Factor (DeepBF), a neural classifier trained on simulated datasets to learn a mapping whose functional constitutes a Bayes factor estimator. The trained mapping enables Bayes factor evaluations over the entire data domain (not just observed data) with negligible additional computational cost. This feature allows us to compute distributions of Bayes factors under each hypothesis and perform model selection inference using tail area evidence measures. Many Bayes factor computation methods are tailored to the observed data [52, 51] which allow only point-evaluations at observed data as opposed to inference using plausible data under each hypothesis. We support our development theoretically by providing sufficient conditions under which our Deep Bayes Factors are consistent estimators of Bayes factors as well as sufficient consistency conditions for model selection. Viewing the Bayes factor through the lens of binary classification aligns with [61], who recast ABC model selection as a classification problem. They employ random forests to select a model by a majority vote. Instead, we focus on binary classification where the purpose is to learn marginal likelihood ratios. Contrary to the method in [61], our strategy circumvents a secondary learning phase for gauging model posterior estimates, delivering results in only one stage. There have several attempts to apply deep learning to the problem of Bayes factor estimation from other perspectives. For example, [39] used normalizing flow approach to estimate the marginal likelihood, and [91] suggested to use f-GAN to improve bridge estimators [1]. While their approaches rely on MCMC sampling from the posteriors to obtain training datasets, our DeepBF is trained solely on simulated data from two competing models, not the posteriors.
@@
Our second contribution is furnishing our methodology with a toolkit for evaluating goodness of model fit as well as goodness of Bayes factors for model selection inference. Our exhaustive evaluation toolkit provides a more comprehensive framework for informed model selection and enables effective objective comparisons of various Bayes factor estimators. Firstly, since model selection cannot be entirely separated from model criticism, we discuss several goodness-of-fit strategies using contrastive learning, which can be used as a prerequisite to Bayes factor evaluations. In particular, we develop a contrastive learning approach to model adequacy assessment based on comparing observed data with data generated from a posterior predictive distribution. We develop a score based on an estimated classifier and gauge model adequacy based its the departure away from $1/2$. While Bayes factor comparisons indicate which of the two considered models is more plausible, our model adequacy framework helps determine whether considered models themselves are compatible with the observed data. Secondly, we evaluate goodness of Bayes factors with visualization tools and validation measures including ROC curves and estimated model prior quality measures. Additionally, while the magnitude of Bayes factors evaluated at observed data traditionally indicates the strength of model evidence [38, 43, 47], recent works complement this by quantifying the amount of evidence using a tail probability [2] derived from simulated data [19, 28, 74, 72]. This improved understanding can lead to calibrations of Bayes factor decision thresholds. Our validation measures, therefore, include the distributional quality of Bayes factor estimates that does not penalize systematic bias or imprecise scale such as area under ROC curve (AUC) and MSE of surprise measures. These evaluations are further enhanced by our rapid and simultaneous computation of our DeepBF estimator. We demonstrate performance of Deep Bayes Factors on several well-studied examples where ABC model choice faces challenges. Lastly, we apply our approach to a real data example on testing for the presence of cognitive biases.
@@ 2 Approximating Bayes Factors via Classification
Our setup consists of an observed data vector $Y^{(n)}_{0}=(Y_{1}^{0},\dots,Y_{n}^{0})\in\mathcal{Y}^{n}$ which could have plausibly arrived from a model $M_{1}$ or a model $M_{2}$. These two models are specified through conditional densities $\pi_{1}(\cdot\,|\>\theta_{1})$ and $\pi_{2}(\cdot\,|\>\theta_{2})$ with unknown model parameters $\theta_{1}\in\Theta_{1}\subseteq\mathbb{R}^{d_{1}}$ and $\theta_{2}\in\Theta_{2}\subseteq\mathbb{R}^{d_{2}}$ that are assigned prior distributions $\pi_{1}(\theta_{1})$ and $\pi_{2}(\theta_{2})$, respectively. The Bayes Factor is then defined as a ratio of the marginal likelihoods evaluated at $Y^{(n)}_{0}$
@@
When likelihoods are analytically unavailable, but can be simulated from, a popular approach to estimating Bayes factors is ABC [27, 84, 16, 66]. We review this approach to build a foundation for our own development.
@@ 2.1 ABC Model Choice
ABC was originally introduced in population genetics [80, 60, 5] for inference about evolution of species. Instead of sampling directly from the posterior, ABC draws candidate parameters from the prior and accepts them if simulated data from the likelihood (given candidate parameters) closely aligns with the observed data. For a comprehensive introduction to ABC we refer to [79]. With some caution, ABC can be also used for model choice.
A common ABC approach to Bayes factor estimation [66] is calculating the ratio of acceptance rates for the two models using ABC reference tables involving simulated data from the marginal likelihood. During the $i^{th}$ of $N$ total ABC iterations, one samples the model index $k^{(i)}\in\{1,2\}$ based on prior model probabilities $\pi(M_{1})$ and $\pi(M_{2})$. Second, parameters $\theta^{(i)}$ are drawn from the prior $\pi_{k^{(i)}}(\cdot)$ and, finally, an $(n\times 1)$ fake data vector $Y^{(i)}$ is simulated from $\pi_{k^{(i)}}(\cdot\,|\>\theta^{(i)})$. The estimate of the Bayes factor based on the ABC lookup table $(k^{(i)},\theta^{(i)},Y^{(i)})_{i=1}^{N}$ is then obtained as
@@
The implementation of $\widehat{BF}^{\rm ABC}_{1,2}(Y_{0}^{(n)})$ in (2) is straightforward and bypasses direct calculations of the likelihood which makes it applicable to a wide range of complex and non-standard models. For Gibbs random fields, [27] demonstrated convergence of (2) to the true Bayes factor. [66] cautioned against inherent bias in ABC-based Bayes factor estimates due to the information loss induced by summary statistics. Their core intuition is that, even if the summary statistic $\eta(\cdot)$ is sufficient for both $\pi_{1}(\cdot\,|\>\theta_{1})$ and $\pi_{2}(\cdot\,|\>\theta_{2})$, it may not be sufficient for $\{(k,\pi_{k}(\cdot\,|\>\theta_{k}))\}_{k=1,2}$, rendering it unsuitable for model selection.
Previous studies in ABC literature have also highlighted the relationship between Bayes factor estimation and the classification problem. Given a set of $K$ competing models $M_{1},\cdots,M_{K}$, [18] estimated the posterior probability $\pi(M_{k}\,|\>Y_{0}^{(n)})$ of each model using multinomial logistic regression trained on an ABC filtered reference table using the model category $k^{(i)}$ as a label and summarized data $\eta(Y^{(i)})$ as predictors. A similar concept is presented in [61] who employed two distinct random forests in two stages. First, a random forest classifier is trained. In the second stage, a separate random forest regression is deployed, targeting the error rate of the initial classifier as an estimator of $\pi(M_{k}\,|\>Y^{(n)}_{0})$. Our method operates in a single stage, where a more flexible classifier is deployed and regarded as a function of the Bayes factor. While we also frame Bayes factor estimation as classification, we do not rely on summary statistics and acceptance thresholds.